### v2

# Grid-Agent: An LLM-Powered Multi-Agent System for Power Grid Control

Abstract
Modern power grids face unprecedented complexity from the increasing penetration of Distributed Energy Resources (DERs), the widespread adoption of Electric Vehicles (EVs), and the growing frequency of extreme weather events. This complexity challenges traditional approaches to grid planning, operation, and management. Traditional rule-based systems and numerical optimization approaches often struggle with the scale, dynamics, and adaptability required by modern power networks. This paper introduces Grid-Agent, an autonomous, AI-driven framework that combines Large Language Models (LLMs) within a multi-agent framework to detect and remediate grid violations. Grid-Agent integrates semantic reasoning with numerical precision through a modular agent architecture: a planning agent generates coordinated action sequences using numerical power flow solvers, while a validation agent evaluates system stability and action effectiveness via sandboxed execution with safety rollbacks. To ensure scalability, Grid-Agent incorporates an adaptive multi-scale network representation that dynamically selects optimal encoding schemes based on network size and complexity. The framework enables coordinated violation resolution through optimizing switch configurations, battery deployment, and load curtailment strategies. Experimental results in standard IEEE and CIGRE test systems (IEEE 69-bus, CIGRE MV, and IEEE 30-bus) demonstrate superior violation mitigation performance. Additionally, the framework’s built-in data collection and learning capabilities enable continuous learning and adaptation to diverse network topologies. The autonomous nature of the framework makes it particularly suitable for modern smart grid applications requiring rapid response to dynamic operating conditions.
Index Terms:
Large language models, multi-agent systems, power grid optimization, violation resolution, smart grids, artificial intelligence, network optimization.
I Introduction
Modern power grids face unprecedented challenges due to increased integration of renewable energy, electric vehicles (EVs), increasing frequency of extreme weather events, and aging infrastructure. These factors demand optimization strategies that are not only intelligent and responsive, but also adaptable to highly dynamic and uncertain operating conditions, especially as grid failures and blackouts have been estimated to cost the U.S. economy over \$150 billion annually [1]. Traditional power system optimization methods are based on fixed rule-based logic and deterministic numerical algorithms, often fall short in their ability to generalize, interpret contextual information, or reason semantically—capabilities that are essential for managing the complexity of real-world grid scenarios.
The emergence of Large Language Models (LLMs) presents transformative opportunities for power system applications by enabling the integration of natural language reasoning with numerical precision. Recent advancements in LLM-based multi-agent systems have demonstrated impressive capabilities in solving complex, multi-faceted problems across various domains [2] [3]. However, their application to mission-critical infrastructures such as electric power grids remains largely unexplored.
This paper introduces Grid-Agent, an autonomous, AI-driven framework conceived as an exploration into how LLMs can be applied to power system control, rather than as a finalized, production-ready solution. The primary goal is to push the boundaries of what LLMs and multi-agent systems might bring to future smart grid operations by demonstrating a novel methodology for autonomous grid management. Consequently, this work prioritizes the exploration of agentic reasoning, coordinated planning, and semantic interpretation over the formal optimality of the resulting control actions. The focus is on demonstrating a paradigm where Grid-Agent can interpret network topology, detect electrical violations, and formulate remediation strategies by integrating semantic understanding with domain-specific analysis tools.
The key contributions of this work include:
  • Multi-Agent architecture. We introduce a multi-agent framework that integrates LLM-based planning with numerical validation, enabling autonomous grid violation detection and remediation without human intervention.
  • Adaptive Multi-scale Representation. A dynamic encoding scheme was developed that automatically adjusts network representation granularity based on system size and complexity, ensuring scalability from micro-grids to large distribution networks.
  • Coordinated Action Optimization. A coordinated violation resolution framework that optimizes multi-violation solutions through topology reconfiguration, battery deployment, and demand response that minimize the number of control actions while maximizing violation resolution effectiveness.
  • Comprehensive Experimental Validation. The framework’s performance was validated on standard IEEE and CIGRE test systems, showing superior violation mitigation.
  • Continuous Learning Capability. A training data generation framework that creates high-quality datasets in multiple formats for LLM fine-tuning to enable the system to learn from operational experience, ensuring its performance improves and adapts over time.
The remainder of this paper is organized as follows: Section II reviews related work in power systems optimization and LLM-based multi-agent systems. Section III formulates the power grid violation resolution problem. Section IV presents the Grid-Agent system architecture and methodology. Section V describes the experimental setup and validation. Section VI discusses results and future works. Section VII concludes the achievements.
II Related Work
II-A Power Systems Optimization
Traditional power system optimization has progressed from manual control to sophisticated automated frameworks. Classical methods, such as Optimal Power Flow (OPF) [4], [5], have been widely applied to objectives like peak shaving, cost/loss minimization, and violation mitigation. While mathematically robust, these approaches often face convergence issues due to the complexity of modern grids, which incorporate high levels of distributed renewable energy, highly variable loads, and numerous discrete controllable assets such as switches and tap changers. These factors result in a mixed-integer, non-convex, nonlinear, and unbalanced AC optimal powerflow problem, which is NP-hard and computationally intensive to solve. Furthermore, their rigid mathematical structure makes it difficult to incorporate heuristic operator knowledge or adapt to unforeseen event types. Furthermore, these approaches often address isolated tasks and lack a unified framework for holistic, system-level optimization and control. This highlights the need for a system that can integrate high-level strategic reasoning with precise numerical validation, a core principle of our proposed Grid-Agent framework.
Recent advances in artificial intelligence (AI) have brought machine learning techniques into power systems, with promising results in tasks such as load forecasting [6], renewable generation prediction [7], and fault detection [8]. However, these approaches generally address isolated tasks and lack a unified framework for holistic, system-level optimization and control.
II-B Large Language Models in Engineering Applications
The application of LLMs to engineering domains has gained significant attention following the success of models like GPT-4 and Claude. Recent work has explored LLM applications in software engineering [9], materials science [10], and system design [11].
In the energy sector, preliminary investigations have examined LLMs for grid analytics [12], energy trading [13], and demand response optimization [14]. However, these applications primarily focus on data analysis and prediction rather than direct system control and optimization. This leaves a critical gap in utilizing LLM reasoning for active, real-time control and optimization in critical infrastructure, which is the primary focus of our work.
II-C Multi-Agent Systems with LLMs
The integration of LLMs with multi-agent systems is an emerging and rapidly advancing area of research. Recent studies have demonstrated the potential of LLM-based multi-agent frameworks in domains such as software engineering [15], automated scientific research [16], and collaborative problem-solving [17]. These systems leverage the reasoning capabilities of LLMs while maintaining structured workflows and state management.
However, the application of LLM-based multi-agent systems to critical infrastructure domains like power grids presents unique challenges including safety requirements, real-time constraints, and the need for explainable decisions in mission-critical environments.
III Preliminaries
This section formally defines the power grid violation resolution problem, the underlying physical laws governing the network, and the discrete and continuous control actions available to the agent.
III-A Power Grid Violation Resolution Problem
We model the power network as a graph G=(N,E){G}=({N},{E}), where ${N}=\{1,2,\dots,n\}$ is the set of buses (nodes) and E{E} is the set of transmission lines and transformers (edges). The state of the grid is defined by the voltage magnitude $V_{i}$ and angle $\delta_{i}$ at each bus $i\in{N}$, along with the active power $P_{ij}$, reactive power $Q_{ij}$, and current flow $I_{ij}$ for each line $(i,j)\in{E}$.
A violation $\mathcal{V}$ occurs when the system state breaches predefined operational limits. We consider three primary categories of violations: voltage violations, thermal violations, and disconnected buses.
A voltage violation $\mathcal{V}_{\text{volt}}$ occurs when a bus voltage magnitude $V_{i}$ falls outside its acceptable range $V_{\min},V_{\max}]$. Formally, the set of buses experiencing voltage violations is defined as
$\mathcal{V}_{\text{volt}}=\{i\in N\mid V_{i}<V_{\min}\text{ or }V_{i}>V_{\max}\}.$ (1)
A thermal violation $\mathcal{V}_{\text{therm}}$ arises when the apparent power flow $S_{ij}$ or current $I_{ij}$ on a line exceeds its maximum rating. This can be expressed as
$\mathcal{V}_{\text{therm}}=\{(i,j)\in E\mid S_{ij}>S_{ij}^{\max}\text{ or }I_{ij}>I_{ij}^{\max}\}.$ (2)
Finally, a disconnected bus violation $\mathcal{V}_{\text{disc}}$ occurs when a bus becomes electrically isolated from the main grid, often resulting in a complete loss of power supply.
III-B Power Flow Constraints
Any valid solution must satisfy the fundamental power flow equations, which represent the physical laws of electricity. For each bus $i\in{N}$, the net active power ($P_{i}$) and reactive power ($Q_{i}$) injections must be balanced. These nonlinear equations connect the bus voltages and angles across the entire network:
$P_{i}=V_{i}\sum_{j\in{N}}V_{j}[G_{ij}\cos(\delta_{i}-\delta_{j})+B_{ij}\sin(\delta_{i}-\delta_{j})]$ (3)
$Q_{i}=V_{i}\sum_{j\in{N}}V_{j}[G_{ij}\sin(\delta_{i}-\delta_{j})-B_{ij}\cos(\delta_{i}-\delta_{j})]$ (4)
where $G_{ij}$ and $B_{ij}$ are the conductance and susceptance terms from the network’s bus admittance matrix, respectively. These equations ensure that any proposed set of control actions results in a physically possible grid state.
III-C Action Space Constraints
The control actions available to Grid-Agent are constrained by the physical and operational capabilities of the grid equipment. The action space consists of switch operations, load curtailment, and battery placement with dispatch.
Switch operations are represented as binary decisions to open or close a switchable line $(i,j)$ within the set of designated switchable lines ${E}_{\text{sw}}\subseteq{E}$. For each switchable line, the state variable is defined as
$s_{ij}\in\{0,1\},\quad\forall(i,j)\in{E}_{\text{sw}}.$ (5)
Load curtailment is modeled as a continuous reduction in the active power demand $P_{i}^{\text{load}}$ at buses belonging to the set of curtailable loads ${N}_{\text{L}}\subseteq{N}$. The reduction is parameterized by a curtailment fraction $\gamma_{i}$, resulting in a new load value given by
$P_{i}^{\text{new\_load}}=P_{i}^{\text{orig\_load}}\cdot(1-\gamma_{i}),$ (6)
$0\leq\gamma_{i}\leq\gamma_{\max},\quad\forall i\in{N}_{\text{L}}.$ (7)
Battery placement and dispatch involve two coupled decisions: a discrete choice of where to place batteries, and a continuous choice of how much active and reactive power to inject or absorb. Placement is modeled as a binary decision variable $b_{i}$ for each bus, constrained by the total number of available batteries $B_{\text{total}}$:
$\sum_{i\in{N}}b_{i}\leq B_{\text{total}},\quad\text{with }b_{i}\in\{0,1\}.$ (8)
Once placed, each battery can dispatch active power $P_{B,i}$ and reactive power $Q_{B,i}$, subject to its maximum power rating $S_{B}^{\max}$. Importantly, dispatch is only allowed if a battery is installed at the corresponding bus ($b_{i}=1$). The coupling between placement and operation is enforced as
$P_{B,i}^{2}+Q_{B,i}^{2}\leq(b_{i}\cdot S_{B}^{\max})^{2},$ (9)
$-P_{B}^{\max}\leq P_{B,i}\leq P_{B}^{\max},$ (10)
$-Q_{B}^{\max}\leq Q_{B,i}\leq Q_{B}^{\max}.$ (11)
This formulation guarantees that batteries are both allocated and dispatched within their physical limits, ensuring they can effectively support grid operations.
III-D Large Language Models and Agentic AI
Solving the violation resolution problem requires more than just numerical optimization; it demands a system that can reason about complex, multi-faceted scenarios, interpret heterogeneous data, and formulate strategic plans. This is where Large Language Models (LLMs) and agentic AI frameworks provide a transformative approach.
III-D1 LLMs as Reasoning Engines for Grid Management
Modern LLMs function as powerful reasoning engines. In the context of power systems, they offer several key advantages over traditional algorithms. First, LLMs bring semantic understanding, as they can process and interpret not just raw numerical data, but also heterogeneous sources of information such as network topology descriptions, operational logs, and maintenance reports. This capability enables a more holistic understanding of grid states. Second, they enable contextual reasoning by incorporating high-level domain knowledge and operational constraints, such as prioritizing critical loads or avoiding switching near vulnerable points, which are difficult to encode in conventional optimization models. Finally, recent advancements allow LLMs to operate in a tool-augmented fashion, delegating complex numerical calculations to validated power flow solvers. This integration grounds their high-level reasoning in the physical realities and precise operational constraints of the grid.
III-D2 Multi-Agent Systems for Decomposed Problem Solving
A single monolithic system struggles to manage the entire scope of violation resolution. Agentic AI addresses this limitation by decomposing the problem into specialized sub-tasks, each handled by a dedicated agent. This approach provides several advantages. Different agents can be designed to develop specialized expertise, excelling at specific roles such as topology analysis, action planning, solution validation, or results summarization. In addition, agents can collaborate iteratively, with a Planner agent proposing an action and a Validator agent simulating its impact to provide feedback. This iterative refinement process allows strategies to improve progressively. Finally, the modular structure offers scalability and interpretability: it adapts to grids of varying complexity while providing transparent and auditable reasoning chains through explicit agent interactions.
The multi-agent system can be formulated as
$W=(A,F,C),$ (12)
where $W$ represents the workflow, $A=\{a_{1},a_{2},\ldots,a_{m}\}$ is the set of agents, $F$ defines the flow relationships between agents, and $C$ specifies conditional branching logic. The workflow state $S_{t}$ at iteration $t$ captures the complete system state, including network configuration, detected violations, executed actions, and intermediate results. State transitions are governed by
$S_{t+1}=\mathcal{T}(S_{t},a_{i},\theta),$ (13)
where $\mathcal{T}$ is the transition function, $a_{i}$ is the executing agent, and $\theta$ represents the agent’s parameters and decision logic. This formulation enables systematic exploration of the action space while maintaining solution quality and computational efficiency through intelligent agent coordination.
IV System Architecture and Methodology
The Grid-Agent framework is an autonomous multi-agent system designed to resolve power grid violations by integrating the semantic reasoning of Large Language Models (LLMs) with the numerical precision of validated power system simulations. The methodology is founded on a state-driven workflow where specialized agents collaborate to plan, execute, and validate control actions in a secure, sandboxed environment. This section details the four core pillars of this methodology: the multi-agent workflow, the LLM-guided action planning process, the adaptive network representation scheme, and the integrated safety mechanisms.
IV-A State-Driven Multi-Agent Workflow
The framework’s operation is governed by a cyclic, state-driven workflow, depicted in Fig. 1, which is orchestrated by five specialized agents. The process begins with the Topology Agent, which initializes the resolution process by parsing the grid’s topology and operational data to establish the initial network state and identify existing violations. This foundational context is then passed to the Planner Agent, the core reasoning engine, which leverages its LLM-based intelligence to formulate a strategic, multi-step action plan designed to cohesively resolve the identified issues. Once a plan is formulated, the Executor Agent acts as the interface to the numerical simulation, translating the Planner’s abstract directives into concrete API calls for a power flow solver on a sandboxed copy of the network. Following execution, the Validator Agent ensures the efficacy and safety of the plan by running a new analysis to assess whether violations were resolved without introducing new instabilities. Based on this outcome, it either approves the changes and terminates the workflow or reverts the changes and signals the need for a new planning cycle, ensuring monotonic progress. Finally, upon successful resolution, the Summarizer Agent facilitates explainability by generating a human-readable summary of the solution and structures the entire process into a data entry for continuous learning datasets.
Figure 1: The architecture diagram of the Grid Agent.
IV-B LLM-Guided Coordinated Action Planning
The central task of finding an optimal action sequence is navigated by the Planner Agent. Unlike traditional optimizers, Grid-Agent uses the LLM as a tool-augmented reasoning engine. This is achieved through a comprehensive system prompt that structures the LLM’s task, defining five key elements:
The core of this process is the Planner Agent, which formulates a multi-step plan (Algorithm 1).The agent’s reasoning is guided by a structured system prompt that specifies:
1) Role Definition: The LLM is instructed to act as an expert power system operator, grounding its responses in domain-specific best practices.
2) State Context: It receives a structured representation of the current network state, including topology, component statuses, and a detailed list of all active violations.
3) Action Space and Constraints: The prompt explicitly defines the set of available actions (e.g., update_switch_status, add_battery, curtail_load) and their associated operational constraints (e.g., maximum number of deployable batteries, available curtailable loads).
4) Strategic Guidance: The LLM is guided by a prioritized action policy that mirrors industry best practices: (1) Topology Reconfiguration, (2) Battery Deployment/Dispatch, and (3) Load Curtailment. This constrains the search space and aligns the LLM’s reasoning with established operational logic.
5) Output Schema: To ensure reliable machine integration, the LLM is required to respond exclusively with a machine-parsable list of tool calls in a structured format (e.g., JSON). This eliminates ambiguous natural language and enables direct interfacing with the Executor Agent.
Algorithm 1 LLM-Based Multi-Agent Power Grid Optimization
1:Network NN, Maximum iterations $T_{max}$, Available actions $\mathcal{A}_{available}$
2:Optimized network state $N^{*}$
[…]
6:while $\neg IsResolved(\mathcal{V})$ and $iter<T_{max}$ do
7:  // Multi-Agent Planning Phase
8:  $\mathcal{C}\leftarrow$ Generate Network representation($N_{work}$, $\mathcal{V}$)
9:  $\mathcal{A}\leftarrow$ LLM planning($\mathcal{C}$, $\mathcal{A}_{available}$)
10:  // Action Execution Phase
11:  for $action\in\mathcal{A}$ do
12:   Execute action(actionaction, $N_{work}$)
13:   $\mathcal{V}_{new}\leftarrow$ Analyze violations($N_{work}$)
14:   if $IsResolved(\mathcal{V}_{new})$ then
15:     break
16:   end if
17:  end for
18:  // Learning and Adaptation
19:  $effectiveness\leftarrow$ Evaluate actions($\mathcal{V}$, $\mathcal{V}_{new}$, $\mathcal{A}$)
20:  Update Knowledge(effectivenesseffectiveness)
21:  $\mathcal{V}\leftarrow\mathcal{V}_{new}$
22:  $iter\leftarrow iter+1$
23:end while
[…]
25:return $N_{work}$, explanationexplanation
IV-C Adaptive Multi-Scale Network Representation
A primary challenge in applying LLMs to large-scale networks is the constraint of the model’s context window. To overcome this, Grid-Agent employs an adaptive network representation scheme that dynamically balances information fidelity with computational tractability. The framework automatically selects the level of detail based on network size and complexity:
Full-Component Detail: For smaller networks, the system provides a complete serialization of every component (buses, lines, loads, generators), offering the LLM comprehensive data for fine-grained analysis.
Semantic Graph Abstraction: For larger networks or scenarios with clustered violations, the system generates a semantic graph representation. This abstraction summarizes healthy sections of the grid and focuses on the electrical relationships between violated components and nearby controllable assets, enabling the LLM to identify coordinated, system-level solutions efficiently.
This multi-scale approach ensures the LLM receives contextually relevant information without being overwhelmed by excessive data, enabling the framework to scale from microgrids to large distribution networks.
IV-D Safety, Validation, and Rollback Mechanisms
Operational safety in a critical infrastructure context is paramount, and Grid-Agent incorporates a multi-layered validation strategy to ensure system stability and reliability. The first layer is a Proactive Execution Check, where the Executor Agent performs a preliminary validation before applying any action in the sandbox. This check uses the power flow solver to ensure the proposed API call is syntactically correct and will not cause an immediate simulation convergence failure, acting as a first line of defense. Following execution, the Validator Agent conducts a Post-Hoc State Assessment, a comprehensive evaluation to verify not only that target violations are resolved but also that no new violations have been introduced elsewhere in the network. Finally, to provide a Monotonic Progress Guarantee, an automated rollback mechanism is triggered if the Validator determines that a plan has failed to improve or has worsened the grid state. This mechanism reverts the sandboxed network to its state before the failed attempt, ensuring the system only progresses towards a verifiably better solution. This defense-in-depth approach, combining proactive checks with comprehensive post-hoc validation and automated rollbacks, provides the essential safety guarantees for deploying an AI-driven system in a mission-critical environment.
IV-E Continuous Learning Framework
A key feature of Grid-Agent is its built-in capability for continuous learning. Upon the successful resolution of a violation scenario, the Summarizer Agent documents the entire process. A dedicated data collection module then compiles a structured training instance containing: (1) the initial network state and violations, (2) the final, successful sequence of actions, and (3) a human-readable explanation of why the solution was effective. This creates a high-quality, domain-specific dataset that can be used to fine-tune the underlying LLM, enabling the system to learn from its operational experience and improve its planning and reasoning capabilities over time.
V Experimental Setup and Validation
To evaluate the performance of the Grid-Agent framework, we designed a comprehensive set of experiments using standardized test networks. This section details the test configurations, performance metrics, baseline models used for comparison, and the results of our evaluation across various Large Language Models (LLMs).
V-A Test Network Configuration
We validated Grid-Agent’s performance on a diverse set of standard IEEE and CIGRE test systems to ensure applicability across networks of varying size, topology, and complexity. The specific test cases were generated by introducing multiple, often interacting, violations to stress-test the agent’s reasoning and coordinated control capabilities.
For the IEEE 30-Bus System, two transmission scenarios were created: Case30 Light, a baseline with 2 voltage and 3 thermal violations, and Case30 Medium, a simpler thermal violation scenario.
To simulate severe distribution-level challenges, we used the CIGRE Medium Voltage (MV) Distribution Network, a 15-bus European benchmark, featuring the CIGRE MV Severe scenario with 14 concurrent violations and the CIGRE MV Disconnected scenario, where a fault isolates 5 buses.
Finally, to evaluate scalability, the larger IEEE 69-Bus Distribution System was tested with scenarios including IEEE 69 Large Loads (18 simultaneous violations), IEEE 69 Medium Loads (a highly complex case with 29 violations), and IEEE 69 Disconnected (a fault resulting in 13 disconnected buses).
An automated script was used to generate these violation scenarios, ensuring repeatable and comprehensive testing across different violation types and severity levels. Table I summarizes the specifications of the benchmark networks employed in this study.
V-B LLM Test Suite Configuration
The core of our study involved comparing the performance of the Grid-Agent framework when powered by different LLMs. We selected six representative models to cover a spectrum of capabilities and sizes: gemini-2.5-pro, gemini-2.5-flash, gemini-2.5-flash-lite, gpt-4.1, gpt-4.1-mini, and gpt-4.1-nano. Each LLM was integrated into the Planner Agent role within the same multi-agent architecture. The performance of these configurations was compared with each other to identify the most effective reasoning engines for this task. To prioritize inference speed, a critical factor for real-time applications, the thinking budget for the gemini models was minimized. The specific configurations for each model are detailed in Table II.
TABLE I: Benchmark Network Specifications
Parameter IEEE Case30 CIGRE MV IEEE 69-bus
Buses 30 14 69
Lines 41 15 68
Loads 20 11 59
Controllable Elements 3 9 8
Network Type Transmission Distribution Distribution
Voltage Level 132-33 kV 20 kV 12.66 kV
TABLE II: Large Language Model Configurations for Benchmarking
Model Version Provider Configuration
gpt-4.1 OpenAI Default Config
gpt-4.1-mini OpenAI Default Config
gpt-4.1-nano OpenAI Default Config
gemini-2.5-pro Google Thinking Budget: 128
gemini-2.5-flash Google Thinking Budget: 0
gemini-2.5-flash-lite Google Thinking Budget: 0
V-C Performance Metrics
System performance was evaluated using a suite of quantitative metrics designed to capture the effectiveness, efficiency, and quality of the violation resolution process. The success rate served as the primary metric, defined as the percentage of violation scenarios that were successfully and fully resolved.
To assess the efficiency of control actions, we measured action efficiency, defined as the average number of violations resolved per control action. This metric reflects the agent’s ability to formulate coordinated and impactful solutions rather than relying on redundant or scattered decisions. Another important aspect was convergence speed, which captures the average number of planning–execution–validation iterations required to achieve complete resolution.
Beyond efficiency and speed, we also considered solution quality. This metric emphasizes not only the resolution of violations but also the manner in which solutions are achieved. High-quality solutions minimize the number of control actions and avoid disruptive measures, such as load curtailment, whenever possible. Finally, runtime was recorded as the average wall-clock time, measured in seconds, required to reach a final solution.
Together, these metrics provide a comprehensive evaluation of the system’s performance, balancing effectiveness, efficiency, practicality, and solution robustness.
V-D Experimental Results and Analysis
The experiments demonstrate that the Grid-Agent framework is highly effective at resolving complex grid violations, with performance varying significantly based on the underlying LLM. Figure.2 provides a comprehensive overview of the key performance metrics across all tested models.
Figure 2: Overall performance comparison across six LLMs. (a) Success rate shows the reliability of each model. (b) Average runtime indicates computational efficiency. (c, d) Average iterations and actions measure the conciseness and quality of the generated solutions.
As shown in Fig.2a, the more capable and efficient models, gemini-2.5-flash and gpt-4.1-mini, achieved a perfect 100% success rate across all scenarios. The flagship models, gemini-2.5-pro and gpt-4.1, also performed robustly with an 100% success rate. The smaller models, gemini-2.5-flash-lite and gpt-4.1-nano, struggled with the more complex scenarios, achieving 63% and 44% success rates, respectively.
There is a clear trade-off between performance and speed. Fig.2b shows that gemini-2.5-flash was by far the fastest model, resolving violations in under 6 seconds on average, whereas the larger gpt-4.1 and gemini-2.5-pro models required over 16 seconds. Figs.2c and 2d show that gemini-2.5-flash also found the most efficient solutions, requiring the fewest iterations and actions, indicating a strong inherent capability for structured, multi-step tool-use reasoning.
Figure 3: Comparison of action efficiency, measured as the average number of violations resolved per control action for successful runs. Higher values indicate more effective, coordinated solutions.
Fig.3 highlights the core contribution of coordinated action planning. The metric measures how many violations, on average, are resolved by a single control action. A higher value suggests that the LLM is better at identifying strategic actions—such as a single switch operation—that can resolve multiple downstream issues simultaneously. The gpt-4.1-nano model, despite its low overall success rate, demonstrated exceptionally high efficiency in the cases it did solve, often finding clever, minimalistic solutions. The top-performing models, gemini-2.5-pro and gemini-2.5-flash, also scored well, resolving nearly three violations per action on average. This significantly outperforms traditional approaches that often tackle violations sequentially, demonstrating the value of the LLM’s holistic reasoning in finding efficient, high-quality solutions.
Beyond overall outcomes, we analyzed the strategic choices made by each LLM. Fig.4 breaks down the action strategies, coordination quality, and overall effectiveness.
Figure 4: Analysis of action strategies. (a) Heatmap showing the proportional usage of each action type. (b) Boxplot of coordination scores, indicating the quality of multi-action plans. (c) Violin plot showing the distribution of total actions required per solution. (d) Scatter plot correlating coordination score with success rate.
The heatmap in Fig.4a reveals distinct strategic preferences. Most models favored topology reconfiguration (update_switch_status) and battery deployment (add_battery), aligning with the prioritized guidance provided in the prompt. This preference for high-impact, non-disruptive actions is a desirable characteristic for grid operations. Notably, gpt-4.1-nano almost entirely avoided curtail_load actions (3% usage), which may explain its lower success rate in scenarios where demand response was essential.
Fig.4b shows the distribution of coordination scores, a metric assessing how well actions within a single step are logically grouped to address related violations. Models like gemini-2.5-pro and gemini-2.5-flash-lite consistently achieved high scores, indicating an ability to formulate synergistic action plans. The violin plot in Fig.4c further explores solution efficiency, showing that most successful models converged in under 10 actions. The wide distribution for gemini-2.5-flash-lite suggests it sometimes resorted to less efficient, sequential corrections.
Crucially, Fig. 4d demonstrates a strong positive correlation between the average coordination score and the final success rate. The top-performing models, gemini-2.5-flash and gpt-4.1-mini, are clustered in the top-right quadrant, confirming that the ability to generate well-coordinated, strategic plans is a key determinant of success in this complex problem domain. This validates that the LLM is not merely guessing but is performing effective, structured reasoning.
To evaluate the framework’s scalability, we analyzed performance across the three network categories, which vary in size and topology (Fig. 5).
Figure 5: Performance analysis across different network types. (a) Success rate heatmap by network category. (b) Average runtime heatmap by network category. (c) Boxplot of action efficiency, highlighting high performance on larger radial systems. (d) Scatter plot of runtime versus network size.
As shown in the heatmaps (Fig.5a and 5b), the IEEE 69-bus system proved to be the most challenging, causing the success rates of weaker models to drop and increasing the runtime for all models. The gemini-2.5-flash model was a notable outlier, maintaining both a 100% success rate and the lowest runtime across all network types. The scatter plot in Fig.5d visualizes this scalability directly, plotting runtime against network elements. The shallow slope for gemini-2.5-flash confirms its superior efficiency and suitability for larger-scale applications. This performance on larger networks is enabled by our adaptive multi-scale representation scheme, which effectively manages the token context size without sacrificing critical information.
Fig.5c reveals a key insight: action efficiency increased dramatically for the larger, radial IEEE 69-bus network. In this system, a single, strategic switch operation can reconfigure power flows to resolve numerous downstream voltage violations simultaneously. The ability of the Grid-Agent to identify and execute these high-impact topological actions is a significant advantage over myopic algorithms and demonstrates its capacity for system-level optimization.
VI Discussion
The experimental results demonstrate that Grid-Agent is a potent framework for resolving complex power grid violations. This section interprets these findings, discusses their implications for power systems operations, and considers the critical aspects of scalability, safety, and reliability.
VI-A Implications for Power Systems Operations
Grid-Agent demonstrates the potential for LLM-based systems to enhance power grid operations through intelligent automation combined with human-interpretable reasoning. The system’s ability to provide clear, step-by-step explanations for its decisions addresses a critical need in power systems operations, where understanding the rationale behind control actions is essential for operator confidence and regulatory compliance.
The coordinated approach to violation resolution, as evidenced by the high action efficiency (Fig.3), offers significant advantages over traditional methods. In scenarios involving renewable energy integration, violations are often interconnected across the network. Grid-Agent’s capacity to identify and execute strategic, high-impact actions—such as a single switch operation that resolves multiple downstream issues—is a marked improvement over myopic algorithms that address violations sequentially. This holistic reasoning is crucial for maintaining stability in increasingly dynamic and complex grids.
VI-B Scalability Considerations
The adaptive multi-scale representation framework directly addresses one of the primary challenges in applying LLMs to large-scale power systems: the limited context window. By dynamically adjusting the level of detail provided to the LLM based on network size and complexity, the system maintains practical computational requirements while preserving solution quality. The strong performance on the IEEE 69-bus system, particularly the superior efficiency of the gemini-2.5-flash model, validates this approach.
Future work should examine scalability to transmission-level networks with thousands of buses. Such large-scale applications will likely exceed the capabilities of a single-agent planner and may require further advances in hierarchical network representation and distributed multi-agent coordination, where regional agents manage local violations and communicate high-level state changes to a central coordinating agent. The integration of retrieval-augmented generation (RAG) will also be explored. Instead of serializing large network sections into the prompt, the LLM would query a specialized vector database containing detailed network topology, component specifications, and a history of successful action sequences for similar faults. This approach would significantly reduce the context size passed to the LLM, leading to faster inference, lower operational costs, and the ability to reason over vast historical and real-time data that would otherwise be computationally prohibitive.
Furthermore, while this exploratory study has focused on the feasibility and strategic reasoning capabilities of Grid-Agent, a crucial direction for future work is the formal analysis of solution optimality. We plan to conduct rigorous benchmarking studies comparing the control actions proposed by Grid-Agent against traditional optimization techniques, particularly mixed-integer nonlinear Optimal Power Flow (OPF) solvers. Such a comparison will be essential to quantify the trade-offs between the LLM’s heuristic, rapid, and interpretable decision-making and the mathematically guaranteed optimality of conventional methods. This will help establish clear performance envelopes for where an agentic approach excels and where traditional optimizers remain indispensable, potentially leading to hybrid systems that leverage the strengths of both.
VI-C Safety and Reliability
The comprehensive validation and rollback mechanisms are cornerstones of the Grid-Agent framework, providing essential safety guarantees for critical infrastructure applications. The system’s ability to explore potential solutions within a secure sandbox, verify their efficacy, and automatically revert failed attempts ensures that the operational grid is never exposed to untested or detrimental actions. This demonstrates the feasibility of AI-driven optimization in safety-critical domains.
However, deployment in a live operational environment would require additional safeguards. These should include robust human-in-the-loop oversight, staged deployment protocols (e.g., operating in an advisory-only mode before granting control), and comprehensive testing under a wider diversity of fault conditions and adversarial scenarios to ensure resilience and build operational trust.
VI-D Integrating Reinforcement Learning for Time-Series Control
While Grid-Agent excels at formulating discrete, strategic actions, a key avenue for future improvement is the integration with Reinforcement Learning (RL) agents for optimizing continuous, time-series control problems. In this paradigm, the LLM-based Planner would act as a high-level orchestrator, identifying the need for a dynamic response (e.g., ”Utilize battery storage to smooth a predicted renewable generation drop over the next hour”).
This high-level directive would then be passed to a specialized RL agent trained to generate an optimal, time-dependent action schedule. This would enable Grid-Agent to support complex, continuous control tasks such as generating a precise State of Charge (SOC) trajectory for battery systems or a dynamic load curtailment schedule over a given time horizon. This hybrid approach would combine the LLM’s semantic reasoning and strategic planning with the RL agent’s proficiency in high-dimensional, sequential decision-making, moving beyond single-shot corrections to proactive, continuous grid management.
VII Conclusion
This paper presented Grid-Agent, a novel multi-agent system that successfully integrates the semantic reasoning of Large Language Models with the numerical precision of power systems optimization. The system demonstrates significant improvements in violation resolution efficiency and provides human-interpretable explanations, which are essential for modern power systems operations.
Key achievements of this work include the development of a modular multi-agent architecture for decomposed problem-solving, an adaptive multi-scale representation framework that enables LLM application to larger networks, and a coordinated action optimization strategy that resolves multiple violations with minimal intervention. The integrated safety mechanisms, including sandboxed execution and automated rollbacks, ensure that all proposed solutions are validated before affecting the system, making the framework suitable for critical infrastructure applications.
The experimental validation on standard IEEE and CIGRE test networks confirms the system’s effectiveness across diverse violation scenarios, with particular strengths in complex, multi-violation situations that challenge traditional optimization approaches. The built-in training data generation framework provides a foundation for continuous learning and adaptation, enabling the system to improve over time. Grid-Agent represents a significant step toward building intelligent, adaptive, and trustworthy power grid management systems that combine the reasoning capabilities of large language models with the rigorous requirements of critical infrastructure.
Acknowledgment
The authors acknowledge the support of the open-source community, particularly the developers of pandapower [18] and LangGraph frameworks that enabled this research.